# Accelerated Convergence of Frank–Wolfe Algorithms with Adaptive Bregman Step-Size Strategy

### v1

@@ 5.2 Local Convergence
Without loss of generality, we set $\sigma=1$. If $\rho\leq L$ does not hold, we can use $h_{1}\leq(\rho+L)D^{2}$ as the initial bound.
Moreover, when $P$ is $\alpha$-strongly convex set and $\phi=\frac{1}{2}\|\cdot\|^{2}$, we also establish local linear convergence. Its proof can be found in Appendix A.2.
6 Numerical Experiments
We compare our algorithm with the following algorithms and use the following notation:
@@ 6 Numerical Experiments
Note that we included OpenAFW only for comparison purposes; there is no established convergence theory for AFW with open-loop strategies. In particular, there are no proper drop steps, and the favorable properties of the away-step Frank-Wolfe algorithm are lost; see [Braun2022-ew]. We use $\beta=0.9$, $\eta=0.9$, $\tau=2$, and $\gamma_{\max}=1$ throughout all numerical experiments.
6.1 Nonnegative Linear Inverse Problems
@@
$\min_{x\in\Delta_{n}}\quad f(x):=d(Ax,b),$ (6.1)
where $d(x,y):=\sum_{i=1}^{m}\left(x_{i}\log\frac{x_{i}}{y_{i}}+y_{i}-x_{i}\right)$ is the Kullback–Leibler divergence (see also Example 2.3). Problem (6.1) is convex, while $\nabla f$ is not Lipschitz continuous on $\mathbb{R}_{+}^{n}$. The pair $(f,\phi)$ is $L$-smad on $\mathbb{R}_{+}^{n}$ with $\phi(x)=\sum_{j=1}^{n}x_{j}\log x_{j}$ and $L\geq\max_{1\leq j\leq n}\sum_{i=1}^{m}a_{ij}$ from [Bauschke2017-hg, Lemma 8].
We compared BregFW with EucFW, ShortFW, OpenFW, and the mirror descent algorithm (MD[Nemirovskij1983-sm]. The subproblem of MD can be solved in closed-form for $\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}=1\}$ by [Beck2017-qc, Example 3.71], and this can be readily extended to $\Delta_{n}$. We used $1000$ as maximum iteration limit and we generated $\tilde{A}$ from an i.i.d. normal distribution and set $a_{ij}=|\tilde{a}_{ij}|/\sum_{i=1}^{m}|\tilde{a}_{ij}|$. We also generated $\tilde{x}$ from an i.i.d. uniform distribution in $[0,1]$ and set the ground truth $x^{*}=0.8\tilde{x}/\sum_{j=1}^{n}\tilde{x}_{j}$ so that $x^{*}\in\operatorname{int}\Delta_{n}$. All components of the initial point $x_{0}$ were $1/n$. For $(m,n)=(100,1000)$, Figure 1 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration (left) and the primal gap per second (right). Table 2 shows average numbers of the primal gap, the FW gap, and computational time over 20 different instances for $(m,n)=(100,1000)$. Here, BregFW outperformed other algorithms, both in iterations and time.
@@
Table 2: Average numbers of the primal gap, the FW gap, and computational time (s) over 20 different instances of nonnegative linear inverse problems for $(m,n)=(100,1000)$.
@@
BregFW 6.963691e-08 1.145520e-05 1.457719e-01
EucFW 3.028696e-07 2.922331e-05 1.077826e-01
ShortFW 4.747044e-05 4.613479e-04 6.251837e-02
OpenFW 4.957628e-07 8.538245e-05 6.503594e-02
MD 2.249368e-06 1.721359e-04 1.888499e-01
6.2 $\ell_{p}$oss Problem
We consider the $\ell_{p}$ loss problem [Kyng2015-uo, Maddison2021-nt] to find $x\in P$ such that $Ax\simeq b$, defined by
$\min_{x\in P}\quad\|Ax-b\|_{p}^{p},$ (6.2)
where $A\in\mathbb{R}^{m\times n}$, $b\in\mathbb{R}^{m}$, and $p>1$. Defining $f(x)=\|Ax-b\|_{p}^{p}$, we have
$\nabla f(x)$ $=p\sum_{i=1}^{m}|\langle a_{i},x\rangle-b_{i}|^{p-1}\operatorname{sgn}(\langle a_{i},x\rangle-b_{i}),$
$\nabla^{2}f(x)$ $=p(p-1)\sum_{i=1}^{m}|\langle a_{i},x\rangle-b_{i}|^{p-2}a_{i}a_{i}^{\mathsf{T}},$
where $a_{i}$ is the $i$th row vector of $A$. Problem (6.2) is convex, but $\nabla f$ is not Lipschitz continuous on $\mathbb{R}^{n}$ when $p\neq 2$. Furthermore, when $p<2$, $f$ is not $\mathcal{C}^{2}$ but $\mathcal{C}^{1}$. Therefore, for $21<p<$ 2, $\nabla f$ is also not Lipschitz continuous over compact sets. Since $f$ is convex, the pair $(f,\phi)$ is 1-smad with $\phi=f$.
We use an $\ell_{2}$ norm ball as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{2}\leq 1\}$. We compared BregFW with EucFW, ShortFW, and OpenFW. We generated $A$ from an i.i.d. normal distribution and normalized it so that $\|a_{i}\|=1$. We also generated $\tilde{x}$ from an i.i.d. normal distribution and set $x^{*}=0.8\tilde{x}/\|\tilde{x}\|$ to ensure $x^{*}\in\operatorname{int}P$. The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$. For $(n,m)=(100,100)$ and $p=1.1$, Figure 2 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration and those gaps per second up to 1000 iterations. Table 3 also shows the average performance over 20 different instances. BregFW outperformed the other algorithms. Since $\nabla f$ is not Lipschitz continuous, ShortFW did not reduce the primal and FW gap.
@@ 6.2 $\ell_{p}$oss Problem
Table 3: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of $\ell_{p}$ loss problems for $(m,n)=(1000,100)$.
@@
BregFW 1.056988e-13 3.764084e-01 5.600876e-01
EucFW 6.341301e-10 8.576574e-01 3.138045e-01
ShortFW 2.465054e+01 4.754413e+01 9.198800e-02
OpenFW 1.698968e-02 4.007941e+00 8.713342e-02
6.3 Phase Retrieval
We are interested in phase retrieval, which involves finding a signal $x\in\mathbb{R}^{n}$ such that
$|\langle a_{i},x\rangle|^{2}\simeq b_{i},\quad i=1,\ldots,m,$
where $a_{i}\in\mathbb{R}^{n}$ describes the model and $b\in\mathbb{R}^{m}$ is a vector of measurements. Phase retrieval has a long history. Patterson studied phase retrieval in X-ray crystallography in 1934 [Patterson1934-sg] and 1944 [Patterson1944-nm]. Phase retrieval arises in many fields of science and engineering, such as image processing [Candes2015-mk], X-ray crystallography [Patterson1934-sg, Patterson1944-nm], and optics [Shechtman2015-ic]. Bolte et al. [Bolte2018-zt] applied the Bregman proximal gradient algorithm to phase retrieval.
In applications, $a_{i}$ and $x$ are often complex vectors. Now, we consider real vectors. To achieve the goal of phase retrieval, we focus on the following nonconvex optimization problem:
$\min_{x\in P}\quad f(x):=\frac{1}{4}\sum_{i=1}^{m}\left(|\langle a_{i},x\rangle|^{2}-b_{i}\right)^{2},$
where $P\subset\mathbb{R}^{n}$ is a compact convex set. We introduce some properties of $f$. Let $\phi$ be defined by
$\phi(x)=\frac{1}{4}\|x\|^{4}+\frac{1}{2}\|x\|^{2},$
which is $\sigma$-strongly convex for $\sigma\leq 1$. For any $L$ satisfying
$L\geq\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right),$ (6.3)
the pair $(f,\phi)$ is $L$-smad on $\mathbb{R}^{n}$ [Bolte2018-zt, Lemma 5.1]. The right-hand side of (6.3) is relatively large in applications, while Takahashi et al. derived a smaller lower bound of $L$ [Takahashi2023-uh, Propositions 5 and 6]. In addition, because we have
$\nabla^{2}f(x)=\sum_{i=1}^{m}\left((3|\langle a_{i},x\rangle|^{2}-b_{i})a_{i}a_{i}^{\mathsf{T}}\right),$
the matrix $\nabla^{2}f(x)+\rho I$ is positive semidefinite for $\rho\geq\sum_{i=1}^{m}\|a_{i}\|^{2}|b_{i}|$, i.e., $f$ is $\rho$-weakly convex. For example, $L=\sum_{i=1}^{m}\left(3\|a_{i}\|^{4}+\|a_{i}\|^{2}|b_{i}|\right)\geq\sum_{i=1}^{m}\|a_{i}\|^{2}|b_{i}|=\rho/\sigma$ holds with $\sigma=1$.
@@ 6.3 Phase Retrieval
We use a $K$-sparse polytope as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{1}\leq K,\|x\|_{\infty}\leq 1\}$. We compared BregFW with EucFW, ShortFW, and OpenFW. The maximum iteration is 1000. We generated $a_{i}$, $i=1,\ldots,m$ from an i.i.d. normal distribution and normalized it to have norm 1. We also generated $x^{*}$ from an i.i.d. uniform distribution in $[0,1]$ and normalized $x^{*}$ to have sum 1. The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$. For $(m,n)=(100,2000)$ and $K=200$, Figure 3 shows the primal and FW gaps per iteration and gaps per second. Because $\nabla f$ is not Lipschitz continuous, ShortFW does not converge, and EucFW performs slowly. The primal gap and the FW gap by the Bregman adaptive step-size strategy are the smallest among these step-size strategies. Table 4 shows the average performance over 20 different instances of $(m,n)=(100,2000)$. We compared BregAFW with EucAFW, ShortAFW, and OpenAFW. In only this setting, we generated $x^{*}$ from an i.i.d. uniform distribution in $[0,1]$ and did not normalize it; that is, $x^{*}$ might be in the face of $P$. Figure 4 shows another setting’s results for $(m,n)=(200,200)$ and $K=110$. Table 5 shows the average performance over 20 different instances for $(m,n)=(200,200)$ and $K=110$. The primal gap by BregAFW is the smallest among these algorithms, while ShortAFW has the smallest value of the FW gap.
Figure 3: Phase retrieval for $(m,n)=(100,2000)$ and $K=200$.
@@
Table 4: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(100,2000)$ and $K=200$.
@@
BregFW 3.307714e-09 4.372800e-06 9.532406e-01
EucFW 1.419221e-04 1.271594e+01 8.291155e-01
ShortFW 5.488362e+04 5.389255e+03 1.822046e-01
OpenFW 4.710460e-08 2.747374e+00 1.945798e-01
Figure 4: Phase retrieval for $(m,n)=(200,200)$ and $K=110$ via away-step FW algorithms.
Table 5: Average numbers of the primal gap, FW gap, and computational time (sec) over 20 different instances of phase retrieval for $(m,n)=(200,200)$ and $K=110$ via away-step FW algorithms.
@@
BregAFW 4.343027e+00 9.757773e-01 2.333602e-01
EucAFW 4.396803e+00 1.221226e+00 3.942057e-01
ShortAFW 4.409245e+00 2.597756e-02 7.353968e-02
OpenAFW 4.409352e+00 7.199201e-02 5.899826e-02
6.4 Low-Rank Minimization

### v3

@@ D.2 Local Sublinear Convergence
D.5 Local Linear Convergence over Uniformly Convex Sets
To this end, we recall the definition of uniformly convex sets.
Definition D.8 ($(\alpha,p)$-uniformly convex set [Braun2022-ew, Definition 2.18][Kerdreux2021-kc, Definition 1.1]).
[…]
$y+\gamma(x-y)+\gamma(1-\gamma)\cdot\alpha\|x-y\|^{p}z\in P.$
Proposition D.9 (Scaling inequality [Braun2022-ew, Proposition 2.19][Kerdreux2021-kc, Lemma 2.1]).
$\frac{\langle u,x-v\rangle}{\|x-v\|^{p}}\geq\alpha\|u\|.$
Theorem D.10 (Local linear convergence over uniformly convex sets).
Suppose that Assumptions 3.1 and 5.1 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ hold and that $P$ is $(\alpha,p)$-uniformly convex set. Let $\nabla f$ be bounded away from 0, i.e., $\|\nabla f(x)\|\geq c>0$ for all $x\in P$. Let $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|$ be the diameter of $P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|x_{t}-v_{t}\|^{2}},1\right\}$. Then, if $\rho<\mu$ and $\rho\leq L$, it holds that
for all $t\geq 1$ where
Proof.
Let $g_{t}:=\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle$ and $h_{t}:=f(x_{t})-f^{*}$. Lemma 3.4 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\nu=1$ is followed by $f(x_{t})-f(x_{t+1})\geq\frac{g_{t}}{2}\gamma_{t}$. Using Proposition D.9, we have
where the third inequality holds from Lemma 3.3 and $\|\nabla f(x)\|\geq c$, and the last inequality holds because of the local quadratic growth property of $f$. The initial bound $h_{1}\leq LD_{\textup{Euc}}^{2}=2LD^{2}$ holds from $\rho\leq L$ and (D.5). For $q=2$, we have $h_{t}-h_{t+1}\geq\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)\min\left\{\frac{\alpha c}{L},1\right\}\cdot h_{t}$, which implies
Thus, we have the claim. For $p>2$, we use Lemma B.11 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c)^{2/p}}{L}$, and $\theta_{0}=1-2/p$ and obtain the claim. ∎
Note that Assumption 3.1 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ holds when $f$ is $L$-smooth over $P$. If $\rho\leq L$ does not hold, we can use the initial bound $h_{1}\leq\frac{\rho+L}{2}D_{\textup{Euc}}^{2}$ from (D.5). In that case, a local linear rate in Theorem D.10 is unchanged. Moreover, we have local linear convergence without assuming $\|\nabla f(x)\|>0$.
Theorem D.11 (Local linear convergence over uniformly convex sets without $\|\nabla f(x)\|>0$).
Suppose that Assumptions 3.1 and 5.1 with $\phi=\frac{1}{2}\|\cdot\|^{2}$ and $\operatorname{int}\operatorname{dom}\phi=\mathbb{R}^{n}$ hold and that $P$ is $(\alpha,p)$-uniformly convex set. Let $D_{\textup{Euc}}:=\sup_{x,y\in P}\|x-y\|$ be the diameter of $P$. Consider the iterates of Algorithm 1 with $\gamma_{t}=\min\left\{\frac{\langle\nabla f(x_{t}),x_{t}-v_{t}\rangle}{L\|x_{t}-v_{t}\|^{2}},1\right\}$. Then, if $\rho<\mu$ and $\rho\leq L$, it holds that
for all $t\geq 1$ where $c=\frac{\sqrt{2\mu}}{\mu-\rho}$ and
Proof.
Using an argument similar to Theorem D.10, we obtain
where the third inequality holds from the PL inequality (B.6) with $c=\frac{\sqrt{2\mu}}{\mu-\rho}$. Therefore, we use Lemma B.11 with $c_{0}=LD_{\textup{Euc}}^{2}$, $c_{1}=\frac{1}{2}\left(1-\frac{\rho}{\mu}\right)$, $c_{2}=c_{1}\cdot\left(1-\frac{\rho}{\mu}\right)^{1-2/p}\frac{(\alpha c^{-1})^{2/p}}{L}$, and $\theta_{0}=1-1/p$ and obtain the claim. ∎
Appendix E Experiments
We use the following notation:
@@ Appendix E Experiments
E.1 Nonnegative Linear Inverse Problems
@@
$\min_{x\in\Delta_{n}}\quad f(x):=d(Ax,b),$ (E.1)
where $d(x,y):=\sum_{i=1}^{m}\left(x_{i}\log\frac{x_{i}}{y_{i}}+y_{i}-x_{i}\right)$ is the Kullback–Leibler divergence. Problem (E.1) is convex, while $\nabla f$ is not Lipschitz continuous on $\mathbb{R}_{+}^{n}$. The pair $(f,\phi)$ is $L$-smad on $\mathbb{R}_{+}^{n}$ with $\phi(x)=\sum_{j=1}^{n}x_{j}\log x_{j}$ and $L\geq\max_{1\leq j\leq n}\sum_{i=1}^{m}a_{ij}$ from [Bauschke2017-hg, Lemma 8].
We compared BregFW with EucFW, ShortFW, OpenFW, and the mirror descent algorithm (MD[Nemirovskij1983-sm]. The subproblem of MD can be solved in closed-form for $\{x\in\mathbb{R}_{+}^{n}\,|\,\sum_{j=1}^{n}x_{j}=1\}$ by [Beck2017-qc, Example 3.71], and this can be readily extended to $\Delta_{n}$. We used $1000$ as maximum iteration limit and we generated $\tilde{A}$ from an i.i.d. normal distribution and set $a_{ij}=|\tilde{a}_{ij}|/\sum_{i=1}^{m}|\tilde{a}_{ij}|$. We also generated $\tilde{x}$ from an i.i.d. uniform distribution in $[0,1]$ and set the ground truth $x^{*}=0.8\tilde{x}/\sum_{j=1}^{n}\tilde{x}_{j}$ so that $x^{*}\in\operatorname{int}\Delta_{n}$ (random seed 1234). All components of the initial point $x_{0}$ were $1/n$. For $(m,n)=(100,1000)$, Figure 5 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration (left) and the primal gap per second (right). Table 2 shows average numbers of the primal gap, the FW gap, and computational time over 20 different instances for $(m,n)=(100,1000)$. Here, BregFW outperformed other algorithms, both in iterations and time.
@@
Table 2: Average numbers of the primal gap, the FW gap, and computational time (s) over 20 different instances of nonnegative linear inverse problems for $(m,n)=(100,1000)$. The results are reported as mean $\pm$ standard deviation.
@@
BregFW 6.9637e-08 $\pm$ 2.0e-08 1.1455e-05 $\pm$ 1.4e-06 1.4577e-01 $\pm$ 4.1e-02
EucFW 3.0287e-07 $\pm$ 9.4e-08 2.9223e-05 $\pm$ 4.1e-06 1.0778e-01 $\pm$ 5.8e-03
ShortFW 4.7470e-05 $\pm$ 3.5e-06 4.6135e-04 $\pm$ 7.2e-05 6.2518e-02 $\pm$ 3.6e-03
OpenFW 4.9576e-07 $\pm$ 6.3e-08 8.5382e-05 $\pm$ 1.3e-05 6.5036e-02 $\pm$ 6.3e-03
MD 2.2494e-06 $\pm$ 3.1e-07 1.7214e-04 $\pm$ 1.9e-05 1.8885e-01 $\pm$ 6.4e-03
E.2 $\ell_{p}$ Loss Problem
@@
Table 3: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of $\ell_{p}$ loss problems for $(m,n)=(1000,100)$. The results are reported as mean $\pm$ standard deviation.
@@
BregFW 1.0589e-13 $\pm$ 8.6e-15 3.7686e-01 $\pm$ 8.6e-03 5.6037e-01 $\pm$ 1.8e-02
EucFW 6.2980e-10 $\pm$ 9.9e-11 8.5940e-01 $\pm$ 1.0e-01 3.1421e-01 $\pm$ 7.4e-03
ShortFW 2.4697e+01 $\pm$ 1.2e+00 4.7621e+01 $\pm$ 8.6e-01 9.2352e-02 $\pm$ 3.7e-03
OpenFW 1.6961e-02 $\pm$ 9.9e-04 4.$0$$\pm$ 1.0e-01 8.7292e-02 $\pm$ 3.5e-03
@@
We use an $\ell_{2}$ norm ball as a constraint, i.e., $P=\{x\in\mathbb{R}^{n}\,|\,\|x\|_{2}\leq 1\}$. We compared BregFW with EucFW, ShortFW, and OpenFW. We generated $A$ from an i.i.d. normal distribution and normalized it so that $\|a_{i}\|=1$. We also generated $\tilde{x}$ from an i.i.d. normal distribution and set $x^{*}=0.8\tilde{x}/\|\tilde{x}\|$ to ensure $x^{*}\in\operatorname{int}P$ (random seed 1234). The initial point $x_{0}$ was generated by computing an extreme point of $P$ that minimizes the linear approximation of $f$. For $(n,m)=(100,100)$ and $p=1.1$, Figure 6 shows the primal gap $f(x_{t})-f^{*}$ and the FW gap $\max_{v\in P}\langle\nabla f(x_{t}),x_{t}-v\rangle$ per iteration and those gaps per second up to 1000 iterations. Table 3 also shows the average performance over 20 different instances. BregFW outperformed the other algorithms. Since $\nabla f$ is not Lipschitz continuous, ShortFW did not reduce the primal and FW gap.
E.3 Phase Retrieval
We show the average performance for the setting of Section 6.2. Tables 4 and 5 show the average performance over 20 different instances of $(m,n)=(1000,10000)$ and $(m,n)=(2000,10000)$, respectively.
We also compared BregAFW with EucAFW, ShortAFW, and OpenAFW. In only this setting, we generated $x^{*}$ from an i.i.d. uniform distribution in $[0,1]$ and did not normalize it (random seed 1234); that is, $x^{*}$ might be in the face of $P$. Figure 7 shows another setting’s results for $(m,n)=(200,200)$ and $K=110$. Table 6 shows the average performance over 20 different instances for $(m,n)=(200,200)$ and $K=110$. The primal gap by BregAFW is the smallest among these algorithms, while ShortAFW has the smallest value of the FW gap.
Figure 7: Phase retrieval for $(m,n)=(200,200)$ and $K=110$ via AFW.
@@ E.3 Phase Retrieval
Table 4: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(1000,10000)$ and $K=2000$. The results are reported as mean $\pm$ standard deviation.
@@
BregFW $-1$$\pm$ 1.1e-12 2.1518e-02 $\pm$ 5.9e-02 4.6990e+01 $\pm$ 1.5e+01
EucFW 1.7170e+00 $\pm$ 3.9e+00 2.4612e+02 $\pm$ 4.5e+02 4.4986e+01 $\pm$ 3.2e+01
ShortFW 3.2739e+05 $\pm$ 1.9e+04 4.2232e+04 $\pm$ 5.0e+02 1.0942e+01 $\pm$ 3.2e-01
OpenFW 5.2496e-10 $\pm$ 3.9e-11 4.3802e+00 $\pm$ 5.5e-02 6.0353e+01 $\pm$ 1.6e+00
Table 5: Average numbers of the primal gap, FW gap, and computational time (s) over 20 different instances of phase retrieval for $(m,n)=(2000,10000)$ and $K=2000$. The results are reported as mean $\pm$ standard deviation.
algorithm primal gap FW gap time
BregFW 6.6743e-12 $\pm$ 5.5e-13 9.5656e-08 $\pm$ 7.9e-09 7.2592e+01 $\pm$ 8.8e+00
EucFW 3.2499e+00 $\pm$ 7.1e+00 3.1304e+02 $\pm$ 4.8e+02 1.4348e+02 $\pm$ 4.3e+01
ShortFW 6.0937e+04 $\pm$ 2.3e+03 2.5461e+04 $\pm$ 2.3e+02 1.2648e+01 $\pm$ 2.4e-01
OpenFW 9.3513e-11 $\pm$ 4.9e-12 2.6628e+00 $\pm$ 1.7e-02 7.1916e+01 $\pm$ 1.3e+00
Table 6: Average numbers of the primal gap, FW gap, and computational time (sec) over 20 different instances of phase retrieval for $(m,n)=(200,200)$ and $K=110$ via away-step FW algorithms. The results are reported as mean $\pm$ standard deviation.
@@
BregAFW 4.3430e+00 $\pm$ 1.7e+00 9.7578e-01 $\pm$ 9.1e-01 2.3336e-01 $\pm$ 6.4e-02
EucAFW 4.3968e+00 $\pm$ 1.7e+00 1.2212e+00 $\pm$ 2.3e+00 3.9421e-01 $\pm$ 1.6e-01
ShortAFW 4.4092e+00 $\pm$ 1.7e+00 2.5978e-02 $\pm$ 7.3e-03 7.3540e-02 $\pm$ 3.5e-02
OpenAFW 4.4094e+00 $\pm$ 1.7e+00 7.1992e-02 $\pm$ 4.9e-02 5.8998e-02 $\pm$ 2.4e-03
E.4 Low-Rank Minimization